OpenStudy (anonymous):

mandy callie and kate are runnin a race of 1500 meters. Mandy wins At the moment Mandy crosses the finish line, she was 300 meters ahead of Callie and 420 ahead of Kate. If Callie and mandy keep running at the same rates as they have so far, how many meters ahead of Kate will Callie be when she crosses?

OpenStudy (phi):

I would assign variable Rc for the Callie's rate, and Rk for Kate's rate Let's say Mandy finishes the race in T minutes. In that same time Callie ran 1500-300= 1200 meters. Kate ran 1500-420= 1080 meters we can set up an equation, using rate * time = distance Rc*T= 1200 or Rc= 1200/T Rk*T= 1080 or Rk= 1080/T Let's say Callie finishes after To minutes. That means Rc*To =1500 (she finished the race) using Rc= 1200/T we find To= 1500*T/1200 = 5T/4 where T is Mandy's time At To time Kate has run Rk*To= 1080/T * 5*T/4= 5*1080/4= 1350 meters can you finish?

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

im trying but i cant seem to figure it out